Show that ∫10 (sin(1/x))/xn 𝑑π‘₯ ; π‘₯ > 0 convergence absolutely, if 𝑛 < 1.


1
Expert's answer
2021-11-04T09:14:49-0400

We want to show that,

"\int_{0}^{1} \frac{(\sin (1 / x))}{x^{n}} d x, x>0 \text { converges }"

Absolutely if "n<1"

Let, "f(x)=\frac{\sin (1 / x)}{x^{n}}, n>0"

0 is the only point of infinite discontinuity and f does net keep the same sign in the interval [0,1].

"\therefore \quad|f(x)| \&=\frac{\left|\sin \left(\frac{1}{x}\right)\right|}{x^{n}}<\frac{1}{x^{n}}"

Also, "\int_{0}^{1} \frac{1}{x^{n}} d x" converges when "n <1" .

Thus, "\int_{0}^{1}\left|\frac{\sin (1 / x)}{x^{n}}\right| d x" converges if and only if"\quad n>0"

Thus, The Integration "\int_{0}^{1}\left|\frac{\sin (1 / x)}{x^{* n}}\right| d x" converges if and only if "n<1"

Or "\int_{0}^{1} \frac{\sin (1 / x)}{x^{n}} d x" converges absolutely if "n<1".